OpenStudy (ktbug3):

will fan and medal!!! plus testimonial 25.7 is what % of 141?

OpenStudy (ktbug3):

@phi

OpenStudy (ktbug3):

@Qwertty123

OpenStudy (ktbug3):

@AloneS @hakunamatata34 @ilovehorses14

OpenStudy (anonymous):

25.7/141 = ?

OpenStudy (bluebeta):

Stop spamming!

OpenStudy (bluebeta):

I just gave you a website that shows you step by step.

OpenStudy (ktbug3):

ok

OpenStudy (anonymous):

the answer is 18.2269504

OpenStudy (c1ndrella78):

100%/x%=141/25.7 (100/x)*x=(141/25.7)*x - we multiply both sides of the equation by x 100=5.48638132296*x - we divide both sides of the equation by (5.48638132296) to get x 100/5.48638132296=x 18.2269503546=x x=18.2269503546 25.7 is 18.2269503546% of 141 or roughly 18 :)

OpenStudy (phi):

25.7 is what % of 141? try to translate it, like this 25.7 = x% of 141 (x is the unknown ) then you should learn x% of 141 means x% times 141 so write it as 25.7 = x*141 divide both sides by 141, like this \[ \frac{25.7}{141}= x \cdot \frac{141}{141}\]

OpenStudy (phi):

you get x= 25.7/141 change that to a decimal, i.e. use a calculator finally, multiply by 100 to change the decimal to a percent